OpenStudy (jango_in_dtown):

How many integers p satisfy the following properties: i) 1<=p<=5000 ii) [sqrt p]=[sqrt(p+125)] where "[]" is the greatest integer function

OpenStudy (jango_in_dtown):

@pooja195

OpenStudy (jango_in_dtown):

@imqwerty

imqwerty (imqwerty):

Think about this- [sqrt p]=[sqrt(p+125)] for this condition to hold (p+125) and p must lie between squares of two consecutive integers

OpenStudy (jango_in_dtown):

correct

OpenStudy (jango_in_dtown):

then

OpenStudy (jango_in_dtown):



OpenStudy (jango_in_dtown):

now (p+125)-p < (n+1)^2-n^2 from the figure

OpenStudy (jango_in_dtown):

so n>62

imqwerty (imqwerty):

Yeah thats correct

OpenStudy (jango_in_dtown):

and now we need an upper bound

imqwerty (imqwerty):

(p+125)-p < (n+1)^2-n^2 Will putting this "<=" instead of "<" work

OpenStudy (jango_in_dtown):

yeah it will work...

OpenStudy (jango_in_dtown):

but it will not give the exact solution

imqwerty (imqwerty):

I gtgz dude :/ i'll check this when i come back

OpenStudy (jango_in_dtown):

ok thanks.. :)...

OpenStudy (jango_in_dtown):

@imqwerty

OpenStudy (jango_in_dtown):

@Loser66 any idea?

OpenStudy (loser66):

\(p\leq 5000\\\sqrt p\leq \sqrt{5000}\\\lfloor\sqrt{p}\rfloor \leq \lfloor \sqrt{5000}\rfloor = 70\)

OpenStudy (loser66):

So min p =4900 then you have p+125= 5025 floor function of them are equal

OpenStudy (jango_in_dtown):

but p+125 may not lie in the same range as p

OpenStudy (loser66):

hey, the condition is p <= 5000, it doesn't say p+125<=5000

OpenStudy (jango_in_dtown):

yeah...

OpenStudy (loser66):

now, max p:

OpenStudy (loser66):

if p =5000, then p+125 = 71.589... then floor of p \(\neq \)floor of p=125 Hence p cannot excess 5000, ok?

OpenStudy (jango_in_dtown):

yes right

OpenStudy (loser66):

I am sorry for deleting it. You need a firmer logic. Give me few minutes to arrange it

OpenStudy (jango_in_dtown):

No problem. :). Take your time

OpenStudy (loser66):

Now, let A be max p, then \(\sqrt{A + 125} <71\) Hence A <4916, that gives us A =4915 hence there are 15 numbers from 4900 to 4915 which satisfy the conditions

OpenStudy (jango_in_dtown):

@Loser66 the answer is 72

OpenStudy (jango_in_dtown):

@Loser66 I did a computer program, and it is showing 72 solutions, 3970-4916

OpenStudy (jango_in_dtown):

@mathmate any idea?

OpenStudy (mathmate):

Well, the situation boils down to k=63 to 70, such that for each k that you have, there are (k+1)^2-(k^2+125) values of p that satisfy the given condition. If you sum the values, that should add up to 72. This is the main idea, and have not check out the details, such as boundary conditions (like where it starts, and where it stops). Please do some work to verify.

OpenStudy (jango_in_dtown):

Any explanation? The answer is correct though

OpenStudy (jango_in_dtown):

@mathmate

OpenStudy (loser66):

:)

OpenStudy (mathmate):

As I mentioned, the algo has to be checked out, which I did, and found that an adjustment is required. It should be the sum of \(((k+1)^2-125) - k^2\) for k equal to 62 to 70, which still adds up to 72. I will let you work out the details, if not already done. If you still prefer to see my reasoning, you just have to ask, but I will give you some time to think about it. Hint: you have done a brute force calculation. Compare the results with this one and the reasoning will be obvious.

OpenStudy (jango_in_dtown):

@mathmate n^2<p<5000 and so n<70.71 i.e. n varies from 63 to 70

OpenStudy (mathmate):

Exactly, that's the upper limit. Did you figure out how the lower limit was obtained?

OpenStudy (jango_in_dtown):

yeah, the numbers p+125 and p must lie between the squares of two consecutive integers

OpenStudy (mathmate):

Exactly. So everything is clear?

OpenStudy (jango_in_dtown):

why is the the total of each such numbers for is (k+1)^2-(k+125) ??

OpenStudy (jango_in_dtown):

@mathmate

OpenStudy (mathmate):

Take k=63. p=k^2=3969, [sqrt(3969)]=63 p+125=4094, [sqrt(4094)]=[63.98]=63, so 3969 is a candidate we can go on until the next perfect square, so for p=3970, [sqrt(3970)]=[sqrt(4095)=63, so 3970 is a candidate. We know that 4096 is not a candidate because it is the next perfect square. Hence the formula: ((k+1)^2-125)-k^2 is the number of such candidates, ex. ((63+1)^2-125)-63^2)=4096-125 - 3969=3971-3969=2 Note: you quoted the outdated (and incorrect) formula I posted. Please read my recent post with the new formula.